# Two transitions in complex eigenvalue statistics: Hermiticity and integrability breaking

### v1

@@ VI Conclusions
Using the NN and NNN distributions, this work shows that the XXZ model with dissipative disorder exhibits two spectral crossovers as a function of disorder strength, $\gamma$. At $\gamma=0$, the model is a pure integrable XXZ spin chain and its spectral statistics (once restricted to an appropriate symmetry sector to avoid exact degeneracies) corresponds to the 1d Poissonian statistics. As the dissipative disorder strength is increased, but still small, the spectrum begins spreading in the complex plane and the statistics, approximately still integrable, are close to Poisson with an effective dimension $D>1$. As the dissipative disorder is increased further, integrability begins to break and both the NN and NNN distributions are well fitted by the 2dCG with $\beta>0$. At $\gamma=2$, the model becomes fully chaotic with NN and NNN spectral statistics close to those of random matrices of class AI, which is in the same non-Hermitian symmetry class as our model. Beyond $\gamma>2$, we observe a crossover from chaos to integrability. The same crossover from chaos to integrability was detected in [roccati2024diagnosing] through real singular value statistics and is now validated also from complex spectral analysis. In a similar manner to the Hermitian case, increasing the dissipative disorder induces a crossover of the spectral statistics from chaotic—described by random matrix symmetry class AI or the 2dCG statistics with $\beta\approx 1.3-1.4$, to integrable—described by the statistics of the 2d Poisson point process (or equivalently by the 2dCG with $\beta=0$). We tracked the crossover by fitting the NN and NNN spectral distributions to the corresponding value of $\beta$ in the 2dCG, obtaining a consistent value for $\beta$ from both fits. We complemented by studying also the complex spectral ratios as a function of the disorder.
@@ Acknowledgements.
We acknowledge funding from the John Templeton Foundation (JTF Grant 62171). FR acknowledges financial support from the Fulbright Research Scholar Program. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the JTF. The numerical simulations presented in this work were in part carried out using the HPC facilities of the University of Luxembourg. GA was partly supported by the Deutsche Forschungsgemeinschaft (DFG) grant SFB 1283/2 2021–317210226 and a Leverhulme Trust Visiting Professorship, grant VP1-2023-007. He is indebted to the School of Mathematics, University of Bristol, where this research was conducted. AC acknowledges the CNLS at Los Alamos National Laboratory, where part of this research was conducted.
@@ Appendix A Random matrix bulk symmetry classes
This appendix briefly recalls what is known about local statistic in non-Hermitian random matrix theory. We focus on the 3 symmetry classes that are conjectured to exhaust the local bulk statistics [hamazaki2020universality]. Representatives for these 3 bulk symmetry classes are class A, AI and AII. We present NN and NNN spacing distributions and compare to the 2dCG, the class AI being presented in Subsection III.3. Our new results include a comparison of the respective NNN with a fit to the 2dCG with the same $\beta$, for class AI in Fig. 4 (right), and for AII in Fig. 15 (right). For illustration, we also compare the NN and NNN distribution of all 3 random matrix classes in Fig. 14 below. Note that, based on pertubation theory, it was shown in [grobe1989universality, haake1991quantum] that for small spacings the repulsion is cubic, $s\sim s^{3}$, irrespectively of the presence or absence of time reversal symmetry.
Figure 14: The NN (left) and NNN (right) spacing distributions of the 3 random matrix symmetry classes. Namely, in order of increasing maxima, class AI (red dashed line), class A (black full line), and class AII (blue dashed-dotted line). These respectively correspond to an increasing $\beta\approx 1.3-1.4$, $\beta=2$, and $\beta\approx 2.5-2.6$.
@@
$J=\Sigma J^{T}\Sigma,\quad P_{N}^{(\rm{AII}^{\dagger})}(J)=C_{N}e^{-\textrm{Tr}(JJ^{\dagger})},$ (25)
with Gaussian distribution of elements, normalisation $C_{N}^{-1}=(\pi/2)^{2N^{2}-N}$, and $\Sigma$ the skew-Hermitian metric
$\Sigma=\begin{pmatrix}0&-i\mathbb{1}_{N\times N}\\ i\mathbb{1}_{N\times N}&0\end{pmatrix}.$ (26)
The NN and NNN spacing distributions have to be determined numerically again and are shown in Fig. 15 as histograms.
Figure 15: Comparison of the NN (left) and NNN (right) spacing distribution for $500$ matrices of size $50002N=$ 5000 of the symmetry class AII (histogram), with the best fit of the 2dCG data for $\beta=2.5$ (red dashed curve) and $\beta=2.6$ (blue curve). The comparison to NNN is new here.
@@
Also shown is a fit of the two respective spacing distributions to $\beta$ in the 2dCG. The corresponding Kolmogorov-Smirnov distances $(d)$ and standard deviations $(\sigma)$ are given in Table 6, with best fits marked in bold. For the NN the value $\beta=2.6$ is slightly favoured, which was also reported in [akemann2022spacing]. For NNN, $\beta=2.6$ again gives the best fit according to the standard deviations, whereas $\beta=2.5$ is best for the Kolmogorov-Smirnov distant. This shows that also the second next eigenvalue of the 2dCG with its locally logarithmic repulsion gives a consistent approximation of this random matrix symmetry class AII. The decreasing quality of the fit for NNN also shows that this is an approximation, and not an exact analytic description.
For comparison with the quality of fits in class A, where we have analytical predictions, we independently ran numerical simulations with matrices of the same size, see Table 2 at the end of Subsection IV.3.

### v2

@@ VI Conclusions
Using the NN and NNN distributions, this work shows that the XXZ model with dissipative disorder exhibits two spectral crossovers as a function of disorder strength $\gamma$. At $\gamma=0$, the model is a pure integrable XXZ spin chain and its spectral statistics (once restricted to an appropriate symmetry sector to avoid exact degeneracies) corresponds to the 1d Poissonian statistics. As the dissipative disorder strength is increased, but still small, the spectrum begins spreading into the complex plane and the statistics, approximately still integrable, are close to Poisson with an effective dimension $D>1$. As the dissipative disorder is increased further, integrability begins to break and both the NN and NNN distributions are well fitted by the 2dCG with $\beta>0$. At $\gamma=2$, the model becomes fully chaotic with NN and NNN spectral statistics close to those of random matrices of class AI, which is in the same non-Hermitian symmetry class as our model. Beyond $\gamma>2$, we observe a crossover from chaos to integrability. The same crossover from chaos to integrability was detected in [roccati2024diagnosing] through real singular value statistics and is now validated also from complex spectral analysis. In a similar manner to the Hermitian case, increasing the dissipative disorder induces a crossover of the spectral statistics from chaotic—described by random matrix symmetry class AI or the 2dCG statistics with $\beta\approx 1.3-1.4$, to integrable—described by the statistics of the 2d Poisson point process (or equivalently by the 2dCG with $\beta=0$). We tracked the crossover by fitting the NN and NNN spectral distributions to the corresponding value of $\beta$ in the 2dCG, obtaining a consistent value for $\beta$ from both fits. We complemented by studying also the complex spectral ratios as a function of the disorder.
@@ Acknowledgements.
We acknowledge funding from the John Templeton Foundation (JTF Grant 62171) and the Luxembourg National Research Fund (FNR, Attract grant 15382998). FR acknowledges financial support from the Fulbright Research Scholar Program. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the JTF. The numerical simulations presented in this work were in part carried out using the HPC facilities of the University of Luxembourg. This work was partly supported by the Deutsche Forschungsgemeinschaft (DFG) grant SFB 1283/2 2021–317210226 (GA, PP) and a Leverhulme Trust Visiting Professorship, grant VP1-2023-007 (GA). GA is indebted to the School of Mathematics, University of Bristol, where this research was conducted. AC acknowledges the CNLS at Los Alamos National Laboratory, where part of this research was conducted.
@@ Appendix A Random matrix bulk symmetry classes
This appendix briefly recalls what is known about local statistic in non-Hermitian random matrix theory. We focus on the 3 symmetry classes that are conjectured to exhaust the local bulk statistics [hamazaki2020universality]. Representatives for these 3 bulk symmetry classes are class A, AI and AII. We present NN and NNN spacing distributions and compare to the 2dCG, the class AI being analysed in Subsection IV.3. Our new results include a comparison of the respective NNN with a fit to the 2dCG with the same $\beta$, for class AI in Fig. 11 (right), and for AII in Fig. A2 (right). For illustration, we also compare the NN and NNN distribution of all 3 random matrix classes in Fig. A1 below. Note that, based on pertubation theory, it was shown in [grobe1989universality, haake1991quantum] that for small spacings the repulsion is cubic, $s\sim s^{3}$, irrespectively of the presence or absence of time reversal symmetry.
Figure A1: The NN (left) and NNN (right) spacing distributions of the 3 random matrix symmetry classes. Namely, in order of increasing maxima, class AI (red dashed line), class A (black full line), and class AII (blue dashed-dotted line). These respectively correspond to an increasing $\beta\approx 1.3-1.4$, $\beta=2$, and $\beta\approx 2.5-2.6$.
@@
$J=\Sigma_{N}J^{T}\Sigma_{N},\quad P_{N}^{(\rm{AII}^{\dagger})}(J)=C_{N}e^{-\textrm{Tr}(JJ^{\dagger})},$ (25)
with Gaussian distribution of elements, normalisation $C_{N}^{-1}=(\pi/2)^{2N^{2}-N}$, and $\Sigma_{N}$ the skew-Hermitian metric
$\Sigma_{N}=\begin{pmatrix}0&-i\mathbb{1}_{N\times N}\\ i\mathbb{1}_{N\times N}&0\end{pmatrix}.$ (26)
The NN and NNN spacing distributions have to be determined numerically again and are shown in Fig. A2 as histograms.
Figure A2: Comparison of the NN (left) and NNN (right) spacing distribution for $500$ matrices of size $50002N=$ 5000 of the symmetry class AII (histogram), with the best fit of the 2dCG data for $\beta=2.5$ (red dashed curve) and $\beta=2.6$ (blue curve). The comparison to NNN is new here.
@@
Also shown is a fit of the two respective spacing distributions to $\beta$ in the 2dCG. The corresponding Kolmogorov-Smirnov distances $(d)$ and standard deviations $(\sigma)$ are given in Table 6, with best fits marked in bold. For the NN the value $\beta=2.6$ is slightly favoured, which was also reported in [akemann2022spacing]. For NNN, $\beta=2.6$ again gives the best fit according to the standard deviations, whereas $\beta=2.5$ is best for the Kolmogorov-Smirnov distant. This shows that also the second next eigenvalue of the 2dCG with its locally logarithmic repulsion gives a consistent approximation of this random matrix symmetry class AII.
For comparison with the quality of fits in class A, where we have analytical predictions, we independently ran numerical simulations with matrices of the same size, see Table 5 at the end of Subsection V.3.